OpenStudy (anonymous):

physics problem please help a swimmer travels from the south end to the north end of a 30m pool in 15s and makes a return trip to the starting position in 19s what is the average velocity round trip?

OpenStudy (anonymous):

1.78947368 m/s

OpenStudy (radar):

Total distance is 30m + 30m or 60m Total time is 15s + 19s or 34 sec. v=d/t

OpenStudy (radar):

I got something less than @cornell

OpenStudy (anonymous):

how can I find the average velocity

OpenStudy (anonymous):

Take the distance divided by time of each and add them together then divide by two

OpenStudy (radar):

Avewrage velocity is total distance divided by total time, please read my earlier posts.

OpenStudy (anonymous):

oh that makes more sense thanks

OpenStudy (anonymous):

Yep. @radar that makes sense, was overthinking it.

OpenStudy (radar):

The problem is asking for the average of the round trip, not for each oneway averaged. At least that is the way I am reading it , never mind we are now in sync.

OpenStudy (radar):

Is this multiple choice, if so what are your options?

OpenStudy (anonymous):

a. 1.60 m/s to the south b. 2.0 m/s to the north c. 1.60 m/s to the north d. 2.0 m/s to the south

OpenStudy (jh3power):

i'll give you a tip,lol.. when the distance is equal use this equation\[v avg= \frac{2v1v2}{v1+v2}\]

OpenStudy (anonymous):

ok thanks

OpenStudy (radar):

You know this problem may be a problem. Velocity differs from speed, in that it is a vector, in other words speed with a direction.. In this case, he is where he started, his "round trip" velocity averaged might be considered zero. The directions were opposite. Now I am confused.